OpenStudy (anonymous):

A triangular region of a community college parking lot was measured. The measure of the first angle of this triangle is double the measure of the second angle. The measure of the third angle is 33 degrees greater then double the measure of the first angle.what are the measurements of the three angles???

OpenStudy (anonymous):

let x = 2nd angle 1st angle is then 2x 3rd angle = 2 x (2x) + 33 three angles in triangle sum to 180 so: x + 2x + 4x + 33 = 180 7x + 33 = 180 7x = 147 x = 21 2x = 42 4x + 33 = 117 if you check these you will see they add to 180 degrees